# Compact quantum algorithms for time-dependent differential equations

### v2

@@ Abstract
Many claims of computational advantages have been made for quantum computing over classical, but they have not been demonstrated for practical problems. Here, we present algorithms for solving time-dependent PDEs, with particular reference to fluid equations. We build on an idea based on linear combination of unitaries to simulate non-unitary, non-Hermitian quantum systems, and generate hybrid quantum-classical algorithms that efficiently perform iterative matrix-vector multiplication and matrix inversion operations. Though these algorithms are end-to-end, they require relatively low-depth quantum circuits and protect quantum advantage, with the best-case asymptotic complexities, which we show are near-optimal. We demonstrate the performance of the algorithms by conducting: (a) fully gate level, state-vector simulations using an in-house, high performance, quantum simulator called QFlowS; (b) experiments on a real quantum device; and (c) noisy simulations using Qiskit Aer. We also provide device specifications such as error-rates (noise) and state sampling (measurement) to accurately perform convergent flow simulations on noisy devices. The results offer evidence that the proposed algorithm is amenable for use on near-term quantum devices.
@@ I Introduction
Simulating nonlinear phenomena is quite arduous. In the case of hydrodynamic turbulence, for instance, the combination of a wide range of scales and the need for fine computational resolution [yeung2020advancing] makes it quite challenging [iyer2021area, buaria2022scaling]. Similar challenges are encountered in other problems such as glassy and molecular dynamics, protein folding and chemical reactions. Simulating such phenomena requires a paradigm shift in computing—and a strong candidate for it is Quantum Computing (QC). QC may surpass classical digital computing in some cases [alexeev2021quantum, awschalom2022roadmap], but it is yet to demonstrate its power in solving practical problems. See, e.g., Refs. [bharadwaj2020quantum, bharadwajintroduction, bharadwaj2024simulating] for a discussion of this point, especially on fluid dynamics. A general point is that most classical systems are nonlinear, whereas quantum algorithms—consisting of quantum gates and quantum circuits—obey laws of quantum mechanics, which are linear and unitary. Reconciling this mismatch makes linearization of the problem inevitable [lin2022koopman, giannakis2022embedding, liu2021efficient], leading to high dimensional linear problems. It is thus necessary to create efficient quantum algorithms to solve high dimensional linear system of equations. Even without linearization, solving nonlinear partial differential equations (PDEs) would still demand, at the least, an efficient way to iteratively operate general non-unitary and non-hermitian matrices. In both cases, one needs a quantum algorithmic framework to perform iterative matrix-vector products and matrix inversion operations efficiently. They are generally of the form
@@ I.1 Related Work
  • Category A – Quantum Direct Numerical Simulations (QDNS) [harrow2009quantum, berry2014high, berry2017quantum, subacsi2019quantum, liu2021efficient, childs2021high, krovi2023improved, an2023linear, fang2023time, bharadwaj2023hybrid, an2023quantum, berry2024quantum, brearley2024quantum, leong2022variational, lubasch2020variational, ingelmann2024two, wright2024noisy, pool2024nonlinear]
@@
QLSA offers complexity guarantees on quantum advantage and are devoid of barren-plateaus and heuristics from optimization. The QLSA-based methods entail solving a linear system of equations derived from the (non)linear governing PDEs. Such a construct is obtained by a combination of discretization methods (e.g., finite difference or finite element methods) and linearization techniques such as Carleman or Koopman embedding techniques, yielding a high dimensional linear system of equations, which is then solved using QLSA. Starting from the Harrow-Hassidim-Lloyd (HHL) algorithm [harrow2009quantum], these sets of algorithms have undergone a considerable evolution [berry2014high, berry2017quantum, subacsi2019quantum, liu2021efficient, childs2021high, krovi2023improved, an2023linear, bharadwaj2023hybrid, fang2023time, an2023quantum, berry2024quantum], by overcoming several challenges that impeded preceding proposals (see [aaronson2015read, bharadwaj2023hybrid]). This has been aided largely by advances made in improving Hamiltonian simulation algorithms [berry2007efficient, berry2014exponential, berry2015simulating, berry2015hamiltonian, low2017optimal, berry2020time, an2021time] as well as the introduction of the concept of Linear Combination of Unitaries [childs2012hamiltonian, childs2017quantum]. The central challenges of QLSA include (i) linear, non-optimal dependence on matrix system parameters such as grid size and number of time steps, matrix sparsity and condition number, and the specified accuracy; (ii) exponentially growing query complexity, requiring multiple copies of the initial state and repeated measurements and amplitude amplification of intermediate and final solution states; (iii) specific constraints on the properties of matrix operators such as hermiticity, unitarity, positive-definiteness and on the range of eigenvalues; (iv) stability requirements for time marching problems, and (v) the depth of quantum circuits due to quantum phase estimation methods and inefficient LCU decomposition. The more recent works, particularly those based on LCU, have addressed some of these aspects, while some [bharadwaj2023hybrid] have also (in the context of fluid dynamics) attempted to make the algorithms end-to-end, by addressing problems of state preparation and post-processing quantum information.
@@
Category C – Algorithms in this category map the original continuum governing PDEs into a Schrödinger or a Schrödinger-like equation. The algorithm presented in [meng2023quantum] maps the Navier-Stokes equations into an analogous hydrodynamic, nonlinear Schrödinger equation by using a Madelung transformation, which is then simulated as a quantum circuit. The upside of this approach stems from its innate quantum mechanical construction, but the overall complexity of the algorithm scales linearly with the problem size and an asymptotic computational advantage remains elusive. A potential alternative is that of Schrödingerisation [jin2022quantum, jin2023quantum, lu2024quantum, hu2024quantum, jin2024quantum], which offers a better computational complexity compared to other works within this category. The approach builds an exact mapping from a classical PDE into a dilated Schrödinger equation having an extra dimension with an effective Hamiltonian. However, in the absence of linearization methods, such mappings are restricted to a specific set of PDEs rather than to general nonlinear PDE’s such as the Navier-Stokes equations [lu2024quantum]. However, this approach tends to be a natural candidate for analog quantum computers than for current gate-based quantum computers, and exhibit a similar computational complexity as some earlier gate-based QLSA proposals. There is still a non-optimal, linear dependence (gate complexity) on the solution accuracy and time [hu2024quantum] for simulating problems that are similar to those considered in this paper.
@@ I.2 Current Contribution
This work presents a set of hybrid, quantum algorithms based on Linear Combination of Unitaries (LCU) for solving time-dependent PDEs, with an application to unsteady flows. We make use of a particular LCU decomposition, which has been used earlier in a different context of non-hermitian open quantum systems [schlimgen2021quantum] and quantum chemistry applications, and transform it into a broader framework of a hybrid quantum-classical algorithm for iterative matrix operations. This is accomplished by introducing six Time Marching Compact Quantum Circuit (TMCQC) algorithms to solve the time-dependent PDEs, admitting both explicit and implicit time stepping schemes. These algorithms require only two, or at most four, controlled unitaries for each LCU decomposition leading to relatively low-depth quantum circuits. We show that, of the six TMCQCs proposed, the gate complexity (of basic two- and one-qubit gates) is the best for (TMCQC-5,6), and is $\mathcal{O}\Big{(}s\epsilon,\log(N_{g}\tau),\textrm{polylog}(\epsilon/\varepsilon_{U},1/\varepsilon_{N}),\log^{-3}((\kappa-1)^{-1})\Big{)}$. Here, the governing PDEs are discretized using a finite difference method with $N_{g}$ as the grid size, $\tau$ as the number of time steps (thus an overall problem size of $\mathcal{O}(N_{g}\tau)$), $s$ and $\kappa$ being the sparsity and condition numbers of the matrix operators, $\epsilon$ a parameter controlling the LCU decomposition and $\varepsilon_{U}$ and $\varepsilon_{N}$ determine the accuracy of different approximations of the algorithm, as discussed in Section IV. We use these algorithms to simulate a linear advection-diffusion problem and use its analytical and classical solutions as reference for estimating the accuracy of the solutions obtained. The simplicity of this problem makes it an ideal candidate for assessing the performance of the quantum algorithm. We show that the results from the proposed algorithm agree qualitatively and quantitatively with analytical results.
@@
The above complexity scaling is near-optimal in all parameters except $s$ and $\epsilon$. However, $s$ is equal to the order of finite difference scheme and therefore contributes only a small constant prefactor and $\epsilon$, as will shown in Section IV.6, may be maintained close to unity, thus adding only modest overhead to the complexity. There is thus the potential for an overall exponential advantage. The complexity (TMCQC-1,2,4) is either linear in $\tau$, or quadratic at worst, while still scaling optimally in other parameters. The qubit complexity scales as $\mathcal{O}(\log(N_{g}\tau))$ at best and $\mathcal{O}(\log(N_{g}\tau(\log(\varepsilon^{-1}_{N}))/\log((\kappa-1)^{-1})))$ at the worst, which is still optimal in both cases and provides an exponential memory advantage. These complexities are without quantum amplitude amplification, whose inclusion gives an improvement up to quadratic. These are significant advances with respect to existing proposals in categories B, C as well as VQA in category A.
@@
The present end-to-end algorithm includes specific Quantum State Preparation and Quantum Post Processing algorithms to conserve quantum advantage, by introducing both explicit and implicit time stepping methods. We provide a fully gate level implementation of state vector simulations using QFlowS—an in-house, high-performance, quantum simulator [bharadwaj2023hybrid, bharadwaj2024qflows]. We also use a real quantum device (IBM Cairo) and study effects of noise using IBM Qiskit Aer platform and propose circuit designs and strategies for accurate, forward time simulations on near-term quantum devices. By choosing suitable problem sizes, our noisy simulation studies of the proposed quantum circuits indicate that the required gate error rates and execution times for convergent simulations are in the ballpark of the state-of-the-art devices and experimental specifications, thus hinting the possibility of retaining quantum advantage. The study also indicates how noise can be used to our advantage.
@@ II Governing equations
where the velocity varies only along y (wall-normal direction). We set $C=U$ (the constant advection velocity in the x-direction), and we put $D$, representing the diffusion coefficient, to unity. The initial condition $u(L/2,0)$ is chosen as the delta function $\delta(x)$ and the system is subject to a periodic boundary condition $u(0,t)=u(L,t)$. Such a setting also admits an analytical solution [ingelmann2024two], making it an ideal test-case to evaluate the accuracy of the quantum solutions. The algorithmic framework developed here is also agnostic to the linearity or otherwise of the PDEs. For instance, if we were to begin by considering the nonlinear case ($C=\mathbf{u}(x,t)$), a preceding linearization such as a Homotopy Analysis and Carlemann or Koopman method [liu2021efficient, giannakis2022embedding, lin2022koopman, bharadwaj2023quantum] can be applied to first obtain an approximate, higher dimensional linear system of equations. Such a system can then be solved using the proposed methods.
@@ III Linear Combination of Unitaries
(a) The first step is to encode the discretized velocity field, $\mathbf{u}(x,t)$, as the amplitudes of an $n_{g}=\log_{2}(N_{g})$ qubit quantum state, $|\Psi\rangle_{u}=\frac{1}{\mathcal{N}}\sum_{j=0}^{N_{g}-1}u_{j}|j\rangle$, where $\mathcal{N}$ is an appropriate normalization constant such that $||~{}|\Psi\rangle_{u}~{}||=1$. The specific structure of such a vector (for a general $b$) would depend on the numerical scheme. This step is performed by a Quantum State Preparation algorithm, which is made of a quantum circuit that efficiently encodes the input values on to qubits using a set of 2 qubit gates. To be able to achieve quantum advantage it is important that this step is done efficiently, with a complexity less than $\mathcal{O}(N_{g})$. It is, in fact, possible to do so if they meet certain sparsity requirements, or have a specific functional form [bharadwaj2023hybrid, vazquez2022enhancing, mozafari2022efficient, grover2002creating] (denoted as operator $R$ from here on). Here, since we use a delta function as our initial condition, such a state can be easily prepared with a single NOT gate, thus lowering the onus on state preparation.
(b) The next step is to apply the finite difference operator $M$ on the prepared state as $M|\Psi\rangle$ (or iteratively as $M^{\tau}|\Psi\rangle$ to march $\tau$ time steps). This is a non-trivial task, since $M$ is neither unitary nor hermitian in general. To make this possible, we invoke the concept of a linear combination of unitary operators. With this, an arbitrary matrix may be approximated by a weighted sum of unitary operators, which themselves can be broken down into fundamental quantum gates. For this, one could use different algorithms proposed earlier [childs2012hamiltonian, zheng2021universal, berry2014exponential, berry2015simulating]. In this work, however, we redesign a strategy that was previously proposed in the context of simulating non-unitary and non-hermitian quantum systems [schlimgen2021quantum, berry2015simulating] into a PDE solver algorithm. This redesigned algorithm requires only 2, or at most 4, controlled unitaries for the linear unitary decomposition, reducing the circuit depths drastically, making the overall algorithm more amenable to use on near-term devices.
@@
However, to accommodate this reduction, the size of input vector would need to be doubled and given by $\hat{b}=[0,b]$, where one half of the dilated vector is padded with zeros. This doubling would require only one additional qubit. From here on, let $K$ represent the number of unitaries corresponding to either two or four unitaries. We now examine how to implement the decomposition as a quantum circuit.
@@
Here, we drop for brevity the normalization constants as may exist, $|\Psi\rangle_{\perp}$ corresponds to an orthogonal subspace that stores the unwanted remainder from the above operation of eq. (9). The post-selected solution subspace is then re-scaled classically by $\mathcal{O}(|||\Psi\rangle_{u}||/\epsilon)=\mathcal{O}(1/\epsilon)$ to obtain the actual solution. However, this procedure also shows that, the circuit for linearly combining the unitaries applies the operator $M$ only probabilistically, and the solution subspace of the quantum state is prepared with a small but finite success probability $p_{\textrm{succ}}$. Since we are interested in applying such a decomposition for $\tau$ time steps, $p_{\textrm{succ}}$ would decay as function of $f(2^{-K},\epsilon^{-1},\tau,|||\Psi\rangle_{\tau}||^{-1})$. The smaller the $p_{\textrm{succ}}$, the larger is the number of repeated circuit simulations (query complexity or shots) required to measure or sample the solution subspace accurately. However, we design the time marching quantum circuits such that the overall query complexity is still kept near optimal, such that the required number of shots does not blow up exponentially. For instance, we use a Richardson extrapolation strategy (outlined in Section V), with which we can do simulations even when $\epsilon\sim\mathcal{O}(1)$.

### v3

@@ Abstract
Many claims of computational advantages have been made for quantum computing over classical, but they have not been demonstrated for practical problems. Here, we present algorithms for solving time-dependent PDEs, with particular reference to fluid equations. We build on an idea based on linear combination of unitaries to simulate non-unitary, non-Hermitian quantum systems, and generate hybrid quantum-classical algorithms that efficiently perform iterative matrix-vector multiplication and matrix inversion operations. These algorithms are end-to-end, with relatively low-depth quantum circuits that demonstrate quantum advantage, with the best-case asymptotic complexities, which we show are near-optimal. We demonstrate the performance of the algorithms by conducting: (a) fully gate level, state-vector simulations using an in-house, high performance, quantum simulator called QFlowS; (b) experiments on a real quantum device; and (c) noisy simulations using Qiskit Aer. We also provide device specifications such as error-rates (noise) and state sampling (measurement) to accurately perform convergent flow simulations on noisy devices. The results offer evidence that the proposed algorithm is amenable for use on near-term quantum devices.
@@ I Introduction
Simulating nonlinear phenomena is quite arduous. In the case of hydrodynamic turbulence, for instance, the combination of a wide range of scales and the need for fine computational resolution [yeung2020advancing] makes it quite challenging [iyer2021area, buaria2022scaling]. Similar challenges are encountered in other problems such as glassy and molecular dynamics, protein folding and chemical reactions. Simulating such phenomena requires a paradigm shift in computing—and a strong candidate for it is Quantum Computing (QC). QC may surpass classical digital computing in some cases [alexeev2021quantum, awschalom2022roadmap], but it is yet to demonstrate its power in solving practical problems. See, e.g., Refs. [bharadwaj2020quantum, bharadwajintroduction, bharadwaj2024simulating] for a discussion of this point, especially on fluid dynamics. A general point is that most classical systems are nonlinear, whereas quantum algorithms—consisting of quantum gates and quantum circuits—obey laws of quantum mechanics, which are linear and unitary. Reconciling this mismatch makes linearization of the problem inevitable [lin2022koopman, giannakis2022embedding, liu2021efficient, costa2023further, gonzalez2024quantum, bharadwaj2023quantum, liu2023efficient], leading to high dimensional linear problems. It is thus necessary to create efficient quantum algorithms to solve high dimensional linear system of equations. Even without linearization, solving nonlinear partial differential equations (PDEs) would still demand, at the least, an efficient way to iteratively operate general non-unitary and non-hermitian matrices that are also endemic to non-hermitian and open quantum systems. In both cases, one needs a quantum algorithmic framework to perform iterative matrix-vector products and matrix inversion operations efficiently. They are generally of the form
@@ I.1 Related Work
  • Category A – Quantum Direct Numerical Simulations (QDNS) [harrow2009quantum, berry2014high, berry2017quantum, subacsi2019quantum, costa2019quantum, liu2021efficient, childs2021high, costa2022optimal, costa2023further, krovi2023improved, liu2023efficient, an2023linear, fang2023time, bharadwaj2023hybrid, an2023quantum, gonzalez2024quantum, dalzell2024shortcut, berry2024quantum, brearley2024quantum, leong2022variational, lubasch2020variational, ingelmann2024two, wright2024noisy, pool2024nonlinear]
@@
QLSA offers complexity guarantees on quantum advantage and are devoid of barren-plateaus and heuristics from optimization. The QLSA-based methods entail solving a linear system of equations derived from the (non)linear governing PDEs. Such a construct is obtained by a combination of discretization methods (e.g., finite difference or finite element methods) and linearization techniques such as Carleman or Koopman embedding techniques, yielding a high dimensional linear system of equations, which is then solved using QLSA. Starting from the Harrow-Hassidim-Lloyd (HHL) algorithm [harrow2009quantum], these sets of algorithms have undergone a considerable evolution [berry2014high, berry2017quantum, subacsi2019quantum, liu2021efficient, childs2021high, krovi2023improved, an2023linear, bharadwaj2023hybrid, fang2023time, an2023quantum, berry2024quantum, costa2022optimal, costa2023further, dalzell2024shortcut, liu2023efficient], by overcoming several challenges that impeded preceding proposals (see [aaronson2015read, bharadwaj2023hybrid]). This has been aided largely by advances made in improving Hamiltonian simulation algorithms [berry2007efficient, berry2014exponential, berry2015simulating, berry2015hamiltonian, low2017optimal, costa2019quantum, berry2020time, an2021time] as well as the introduction of the concept of Linear Combination of Unitaries (LCU) [childs2012hamiltonian, childs2017quantum]. The central challenges of QLSA include (i) linear, non-optimal dependence on matrix system parameters such as grid size and number of time steps, matrix sparsity and condition number, and the specified accuracy; (ii) exponentially growing query complexity, requiring multiple copies of the initial state and repeated measurements and amplitude amplification of intermediate and final solution states; (iii) specific constraints on the properties of matrix operators such as hermiticity, unitarity, positive-definiteness and on the range of eigenvalues; (iv) stability requirements for time marching problems, and (v) the depth of quantum circuits due to quantum phase estimation methods and inefficient LCU decomposition. The more recent works, particularly those based on LCU, have addressed some of these aspects, while some [bharadwaj2023hybrid] have also (in the context of fluid dynamics) attempted to make the algorithms end-to-end, by addressing problems of state preparation and post-processing quantum information.
@@
Category C – Algorithms in this category map the original continuum governing PDEs into a Schrödinger or a Schrödinger-like equation. The algorithm presented in [meng2023quantum] maps the Navier-Stokes equations into an analogous hydrodynamic, nonlinear Schrödinger equation by using a Madelung transformation, which is then simulated as a quantum circuit. The upside of this approach stems from its innate quantum mechanical construction, but the overall complexity of the algorithm scales linearly with the problem size and an asymptotic computational advantage remains elusive. A potential alternative is that of Schrödingerisation [jin2022quantum, jin2023quantum, lu2024quantum, hu2024quantum, jin2024quantum], which offers a better computational complexity compared to other works within this category. The approach builds an exact mapping from a classical PDE into a dilated Schrödinger equation having an extra dimension with an effective Hamiltonian. However, in the absence of linearization methods, such mappings are restricted to a specific set of PDEs rather than to general nonlinear PDE’s such as the Navier-Stokes equations [lu2024quantum]. However, this approach tends to be a natural candidate for analog quantum computers rather than for current gate-based quantum computers, and exhibit a similar computational complexity as some earlier gate-based QLSA proposals. There is still a non-optimal, linear dependence (gate complexity) on the solution accuracy and time [hu2024quantum] for simulating problems that are similar to those considered in this paper.
@@ I.2 Current Contribution
This work presents a set of hybrid, quantum algorithms based on LCU for solving time-dependent PDEs, with an application to unsteady flows. We make use of a particular LCU decomposition, which has been used earlier in a different context of non-hermitian open quantum systems [schlimgen2021quantum] and quantum chemistry applications, and transform it into a broader framework of a hybrid quantum-classical algorithm for iterative matrix operations. This is accomplished by introducing six Time Marching Compact Quantum Circuit (TMCQC) algorithms to solve the time-dependent PDEs, admitting both explicit and implicit time stepping schemes. These algorithms require only two, or at most four, controlled unitaries for each LCU decomposition leading to relatively low-depth quantum circuits. We show that, of the six TMCQCs proposed, the gate complexity (of basic two- and one-qubit gates) is the best for (TMCQC-5,6), and is $\mathcal{O}\Big{(}s\epsilon,\log(N_{g}\tau),\textrm{polylog}(\epsilon/\varepsilon_{U},1/\varepsilon_{N}),\log^{-3}((\kappa-1)^{-1})\Big{)}$. Here, the governing PDEs are discretized using a finite difference method with $N_{g}$ as the grid size, $\tau$ as the number of time steps (thus an overall problem size of $\mathcal{O}(N_{g}\tau)$), $s$ and $\kappa$ being the sparsity and condition numbers of the matrix operators, $\epsilon$ a parameter controlling the LCU decomposition and $\varepsilon_{U}$ and $\varepsilon_{N}$ determine the accuracy of different approximations of the algorithm, as discussed in Section IV. We use these algorithms to simulate a linear advection-diffusion problem and use its analytical and classical solutions as reference for estimating the accuracy of the solutions obtained. The simplicity of this problem makes it an ideal candidate for assessing the performance of the quantum algorithm. We show that the results from the proposed algorithm agree qualitatively and quantitatively with analytical results.
@@
The above complexity scaling is near-optimal in all parameters except $s$ and $\epsilon$. By optimal we mean that the algorithmic complexity (qubit, query and two-qubit gate complexities) scale logarithmically or at most poly-logarithmically in the system parameters. By near-optimal, we imply that complexity is optimal in most parameters but a few, in this case $s$ and $\epsilon$. However, $s$ is equal to the order of finite difference scheme and therefore contributes only a small constant prefactor and $\epsilon$, as will shown in Section IV.6, may be maintained close to unity, thus adding only modest overhead to the complexity. There is thus the potential for an overall exponential advantage. The complexity (TMCQC-1,2,4) is either linear in $\tau$, or quadratic at worst, while still scaling optimally in other parameters. The qubit complexity scales as $\mathcal{O}(\log(N_{g}\tau))$ at best and $\mathcal{O}(\log(N_{g}\tau(\log(\varepsilon^{-1}_{N}))/\log((\kappa-1)^{-1})))$ at the worst, which is still optimal in both cases and provides an exponential memory advantage. These complexities are without quantum amplitude amplification, whose inclusion gives an improvement up to quadratic. These are significant advances with respect to existing proposals in categories B, C as well as VQA in category A.
@@
The present end-to-end algorithm includes specific Quantum State Preparation and Quantum Post Processing algorithms to to demonstrate quantum advantage, by introducing both explicit and implicit time stepping methods. We provide a fully gate level implementation of state vector simulations using QFlowS—an in-house, high-performance, quantum simulator [bharadwaj2023hybrid, bharadwaj2024qflows]. We also use a real quantum device (IBM Cairo) and study effects of noise using IBM Qiskit Aer platform and propose circuit designs and strategies for accurate, forward time simulations on near-term quantum devices. By choosing suitable problem sizes, our noisy simulation studies of the proposed quantum circuits indicate that the required gate error rates and execution times for convergent simulations are in the ballpark of the state-of-the-art devices and experimental specifications, thus hinting the possibility of retaining quantum advantage. The study also indicates how noise can be used to our advantage.
@@ II Governing equations
where the velocity varies only along y (wall-normal direction). We set $C=U$ (the constant advection velocity in the x-direction), and we put $D$, representing the diffusion coefficient, to unity. The initial condition $u(L/2,0)$ is chosen as the delta function $\delta(x)$ and the system is subject to a periodic boundary condition $u(0,t)=u(L,t)$. Such a setting also admits an analytical solution [ingelmann2024two], making it an ideal test-case to evaluate the accuracy of the quantum solutions. The algorithmic framework developed here is also agnostic to the linearity or otherwise of the PDEs. For instance, if we were to begin by considering the nonlinear case ($C=\mathbf{u}(x,t)$), a preceding linearization such as a Homotopy Analysis and Carlemann or Koopman method [liu2021efficient, liu2023efficient, costa2023further, giannakis2022embedding, lin2022koopman, bharadwaj2023quantum, gonzalez2024quantum] can be applied to first obtain an approximate, higher dimensional linear system of equations. Such a system can then be solved using the proposed methods.
@@ III Linear Combination of Unitaries
(a) The first step is to encode the discretized velocity field, $\mathbf{u}(x,t)$, as the amplitudes of an $n_{g}=\log_{2}(N_{g})$ qubit quantum state, $|\Psi\rangle_{u}=\frac{1}{\mathcal{N}}\sum_{j=0}^{N_{g}-1}u_{j}|j\rangle$, where $\mathcal{N}$ is an appropriate normalization constant such that $||~{}|\Psi\rangle_{u}~{}||=1$. The specific structure of such a vector (for a general $b$) would depend on the numerical scheme. This step is performed by a Quantum State Preparation algorithm, which is made of a quantum circuit that efficiently encodes the input values on to qubits using a set of 2 qubit gates. To be able to achieve quantum advantage it is important that this step is done efficiently, with a complexity less than $\mathcal{O}(N_{g})$. It is, in fact, possible to do so if they meet certain sparsity requirements, or have a specific functional form [bharadwaj2023hybrid, vazquez2022enhancing, mozafari2022efficient, grover2002creating] (denoted as operator $R$ from here on). Here, since we use a delta function as our initial condition, such a state can be easily prepared with a single NOT gate, thus lowering the onus on state preparation. However, as noted in [bharadwaj2023hybrid], while arbitrary states can be exponentially expensive to prepare, certain generic states (more general than a delta function) with specific features such as convex functional forms, sparsity and random initializations can be prepared efficiently. Furthermore, problems in fluid dynamics such as turbulence are flexible enough to admit a wide range of initial conditions including random states, thus alleviating the state preparation overhead to a certain extent.
(b) The next step is to apply the finite difference operator $M$ on the prepared state as $M|\Psi\rangle$ (or iteratively as $M^{\tau}|\Psi\rangle$ to march $\tau$ time steps). This is a non-trivial task, since $M$ is neither unitary nor hermitian in general. To make this possible, we invoke the concept of a linear combination of unitary operators. With this, an arbitrary matrix may be approximated by a weighted sum of unitary operators, which themselves can be broken down into fundamental quantum gates. For this, one could use different algorithms proposed earlier [childs2012hamiltonian, zheng2021universal, berry2014exponential, berry2015simulating]. In this work, however, we redesign a strategy that was previously proposed in the context of simulating non-unitary and non-hermitian quantum systems [schlimgen2021quantum, berry2015simulating] into a PDE solver algorithm. This redesigned algorithm requires only 2, or at most 4, controlled unitaries for the linear unitary decomposition, reducing the circuit depths drastically, making the overall algorithm more amenable to use on near-term devices. Further, invoking specific techniques that we explain in Section V, these unitaries would then correspond to 15 and 139 two-qubit gates for two and four qubit circuits, with depths of 10 and 95 gates, respectively.
@@
Where, $e^{-i\epsilon\hat{M}}=\hat{U}_{0}$ and $-e^{i\epsilon\hat{M}}=\hat{U}_{1}$. However, to accommodate this reduction, the size of input vector would need to be doubled and given by $\hat{b}=[0,b]$, where one half of the dilated vector is padded with zeros. This doubling would require only one additional qubit. From here on, let $K$ represent the number of unitaries corresponding to either two or four unitaries. We now examine how to implement the decomposition as a quantum circuit.
@@
For brevity, we drop the required normalization constants and $|\Psi\rangle_{\perp}$ corresponds to an orthogonal subspace that stores the unwanted remainder from the above operation of eq. (9). The post-selected solution subspace is then re-scaled classically by $\mathcal{O}(|||\Psi\rangle_{u}||/\epsilon)=\mathcal{O}(1/\epsilon)$ to obtain the actual solution. However, this procedure also shows that, the circuit for linearly combining the unitaries applies the operator $M$ only probabilistically, and the solution subspace of the quantum state is prepared with a small but finite success probability $p_{\textrm{succ}}$. Since we are interested in applying such a decomposition for $\tau$ time steps, $p_{\textrm{succ}}$ would decay as function of $f(2^{-K},\epsilon^{-1},\tau,|||\Psi\rangle_{\tau}||^{-1})$. The smaller the $p_{\textrm{succ}}$, the larger is the number of repeated circuit simulations (query complexity or shots) required to measure or sample the solution subspace accurately. However, we design the time marching quantum circuits such that the overall query complexity is still kept near optimal, such that the required number of shots does not blow up exponentially. For instance, we use a Richardson extrapolation strategy (outlined in Section V), with which we can do simulations even when $\epsilon\sim\mathcal{O}(1)$.